?? integration by parts

Grade:12th pass

3 Answers

grenade
8 years ago
A rational function can be integrated by expressing it as the sum of partial fractions.
gdbusher
8 years ago
will you please explore
ng29
8 years ago

 integration by parts is a theorem that relates the integral of a product of functions to the integral of their derivative and antiderivative. It is frequently used to transform the antiderivative of a product of functions into an antiderivative for which a solution can be more easily found. The rule can be derived in one line simply by integrating the product rule of differenttiation

 

If u = u(x) and du = u′(xdx, while v = v(x) and dv = v′(xdx, then integration by parts states that:

APPROVE IF USEFUL     \int u(x) v`(x) \, dx = u(x) v(x) - \int v(x) \, u`(x)  dx